Use the order completeness property to show that the set
S={n/(n+1): n∈N}

has a
supremum and infimum.
1
Expert's answer
2020-03-06T11:17:57-0500

1 STEP: Search for the supremum of a given sequence



"\frac{n}{n+1}=\frac{(n+1)-1}{n+1}=1-\frac{1}{n+1}"



Conclusion,


"S_n=1-\frac{1}{n+1}<1,\quad\forall n\in\mathbb{N}\\[0.5cm]
\boxed{\sup(S_n)=1}"



2 STEP: Search for the infimum of a given sequence



"1+n\le n+n,\forall n\in\mathbb{N}\longrightarrow\\[0.5cm]
\frac{1}{1+n}\ge\frac{1}{n+n}\longrightarrow\frac{n}{n+1}\ge\frac{n}{n+n}=\frac{1}{2}"

Conclusion,


"S_n\ge\frac{1}{2},\quad\forall n\in\mathbb{N}\\[0.5cm]
\boxed{\inf(S_n)=\frac{1}{2}}"